OpenStudy (anonymous):

probability mass function? I bought 12 eggs which weight 54,55,57,59,59,61,63,66,66,67,67,and 77 grams. let x be the weight of a randomly selected egg. what is the PMF of x?

OpenStudy (tkhunny):

Are these 12 eggs suppose to represent th ENTIRE population of eggs? 5 4 5 5 5 7 5 99 6 1 6 3 6 66 6 77 7 7 Looks like a bunch of p() = 1/12 and a few p() = 1/6

OpenStudy (anonymous):

i think so. Im not completely confident what im even required to do for this question. am i suppose to get an equation or what?

OpenStudy (tkhunny):

Just list the probability masses: p(54) = 1/12 p(55) = 1/12 p(57) = 1/12 p(59) = 1/6 etc... Make sure it addes up to unity (1) when you are done.

OpenStudy (anonymous):

and is that the entire PMF?

OpenStudy (anonymous):



OpenStudy (anonymous):

like that? but continue for the rest im assuming

OpenStudy (tkhunny):

...all the way to 77.

OpenStudy (anonymous):

alright thank you